Solve by graphing
3x-2y=12
7x+2y=7

Answer by ewatrrr(24785)


Hi
Solve by graphing
3x-2y=12 (0,-6) (4,0) (2,-3)
7x+2y=7 (0, 7/2) (1,0) (-1,7)
Pt(19/10, -63/20) is the solution for the system (~2,~(-3)) an approxiamation